if 2 1/2 kg of beef is sold for $40.00 calculate at this time the same rate,the cost of 6 kg of beef

Answer by checkley79(3341)

2.5/40=6/x
2.5x=6*40
2.5x=240
x=240/2.5
x=$96.00 for the 6 kg of beef.
Proof:
2.5/40=6/96
240=2.5*96
240=240